OpenStudy (anonymous):

What is the axis of symmetry and the x and y intercept of f(x)=x^2-4x

OpenStudy (amistre64):

y int = 0 becasue there is no constant on the end of your equation

OpenStudy (amistre64):

x(x-4) = = gets your x ints x = 0 and x = 4

OpenStudy (amistre64):

the axis of symmetry is where your vertex is; at -b/2a 4/2(1) = 2 x = 2 is your axis of symmetry

OpenStudy (anonymous):

ok, so whenever there is no constant the y intercept is always 0

OpenStudy (amistre64):

you could have also just wnet half way between the xints :)

OpenStudy (amistre64):

that last number tells you where it crosses the y axis; if there is no number that just means that it is actually a +0 that they decide not to write... x^2 -4x = x^2 -4x +0 right?

OpenStudy (anonymous):

right! I have the vertex as 2,-4 is that correct?

OpenStudy (amistre64):

when you plug in the x=2, you get the answer for the "y" part, so yes :)

OpenStudy (anonymous):

Ugh!! Thanks! I am going crazy here with this math class!!

OpenStudy (amistre64):

i was bored to tears in college algebra..... and in trig..... and probably will be this summer in stats and calculus1

OpenStudy (anonymous):

I usually do well in math but I am taking on line classes so not having the teacher there to explain things is what is making it difficult!

OpenStudy (amistre64):

stats should be ok....i aint to brushed up on it :)

OpenStudy (anonymous):

I got a D in statistics!

OpenStudy (amistre64):

..... well...... if you did your best, then that is a mighty fine D :)

OpenStudy (anonymous):

I think it was the teacher, he was all over the place! I am a nurse and always was very good in math.

OpenStudy (amistre64):

rn lpn..?? my mother was a lpn and she loved it, grandma and aunt and family are nurses

OpenStudy (anonymous):

I have been an LPN for 12 years and i love it but i am going back to school to be a teacher

OpenStudy (amistre64):

after 20 years of the workplace, I decided to go back to school as well; getting me a bachelors in business and a masters in math....end goal is to be a college math teacher...

OpenStudy (anonymous):

oh wow, good for you! I am leaving nursing for my kids, I love being able to spend time with them, they are 6 and 11 and growing so fast!

OpenStudy (amistre64):

yeah, mine are 13 and 15 now :)

OpenStudy (anonymous):

They grow up too fast!!

OpenStudy (amistre64):

:) and too loud lol

OpenStudy (anonymous):

oh yeah! My 11 yr old has quite the little attitude as well!